Convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3".
a) 16 3/2 = 64
b) ex = 5
B. Convert to exponential form. For example, the exponential form of "log5 25 = 2" is "52 = 25".
a) log 3 27 = 3
b) log e 1 = 0
c) log 125 25 = 2/3

Answer by stanbon(75887)

Note: Both exponential and logarithm equations have three parts:
exponent or log
base
result or inverse log
---------------------------------

Convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3".
a) 16 3/2 = 64
exponent or log is (3/2)
base is 16
result or inverse log is 64
Rewrite as
log(base 16) 64 = (3/2)
------------------------------
b) e^x = 5
exponent of log is "x"
base is e
result of inverse log is 5
Rewrite as
log(base e) 5 = x
-------------------------
B. Convert to exponential form. For example, the exponential form of "log5 25 = 2" is "52 = 25".
a) log 3 27 = 3
log or exponent is 3
base is 3
result of inverse log is 27
Rewrite as 3^3 = 27
-------------------------
b) log e 1 = 0
exponent or log is 0
base is e
result or inverse log is 1
Rewrite as e^0 = 1
==============================
c) log 125 25 = 2/3
exponent or log is 2/3
Base = 125
Result or inverse log is 25
Rewrite as 125^(2/3) = 25
======================
Cheers,
Stan H.